An investment project has annual cash inflows of $5,800, $6,900, $7,700 for the next four years, respectively, and $9,000, and a discount rate of 14 percent. What is the discounted payback period for these cash flows if the initial cost is $9,000?

Multiple Choice 3.48 years

2.49 years

1.24 years

0.74 years

1.74 years
Answer #1

Present value of year 1 cash flow = 5800 / (1 + 0.14) = 5,087.7193

Present value of year 2 cash flow = 6,900 / (1 + 0.14)2 = 5,309.3259

Present value of year 3 cash flow = 7,700 / (1 + 0.14)3 = 5,197.28068

Present value of year 4 cash flow = 9,000 / (1 + 0.14)4 = 5,328.7225

Cumulative cash flow for year 0 = -9000

Cumulative cash flow for year 1 = -9000 + 5,087.7193 = -3,912.2807

Cumulative cash flow for year 2 = -3,912.2807 + 5,309.3259 = 1,397

3,912.2807 / 5,309.3259 = 0.74

Discounted Payback period = 1 + 0.74 = 1.74 years